# references.bib

### v3

@@ 2.2. Parabolic induction
For $\lambda\in X_{*}(T)$, we let $\mathbf{Q}_{V^{\lambda}/G^{\lambda}}^{\mathrm{vir}}\coloneqq\mathbf{Q}_{V^{\lambda}/G^{\lambda}}\otimes\mathscr{L}^{-d_{\lambda}/2}=\mathcal{IC}(V^{\lambda}/G^{\lambda})$ be the intersection complex monodromic mixed Hodge module of the smooth stack $V^{\lambda}/G^{\lambda}$. We refer to [davison2020cohomological] for the necessary background regarding monodromic mixed Hodge modules. Alternatively, for the purposes of this paper, the reader may prefer to consider instead constructible sheaves. The Tate twist $-\otimes\mathscr{L}$ is then replaced by the shift $[-2]$.
@@
Moreover, the map $p_{\lambda}$ is proper (Lemma 2.1), and so by dualizing the adjunction map
@@
by composing $(\imath_{\lambda})_{*}(\pi_{\lambda})_{*}$ applied to (2.1) with $\pi_{*}$ applied to (2.2) twisted by $\mathscr{L}^{r_{\lambda}+d_{\lambda}/2}$.
@@
Let $G$ be a symmetric representation of $G$. The induction morphism $\mathrm{Ind}_{\lambda}\colon\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/G^{\lambda})\rightarrow\mathrm{H}^{*+d}(V/G)$ preserves cohomological degrees.
@@
This is straightforward since the induction at the level of cohomology vector spaces is given by taking the derived global section of the morphism of complexes (2.3) (the sheafified induction) and by Lemma 2.4, $d/2=r_{\lambda}+d_{\lambda}/2$ and so in the symmetric case, the right-hand-side of (2.3) is precisely $\pi_{*}\mathbb{D}\mathbf{Q}_{V/G}^{\mathrm{vir}}$. Moreover, $\mathbb{D}\mathbf{Q}_{V/G}^{\mathrm{vir}}\cong\mathbf{Q}_{V/G}^{\mathrm{vir}}$ and $\mathbb{D}\mathbf{Q}_{V^{\lambda}/G^{\lambda}}^{\mathrm{vir}}\cong\mathbf{Q}_{V^{\lambda}/G^{\lambda}}^{\mathrm{vir}}$ since $V/G$ and $V^{\lambda}/G^{\lambda}$ are smooth and the chosen shifts make the considered constant sheaves Verdier self-dual. ∎
@@ 2.3. Augmentation
Let $\lambda\in X_{*}(T)$. We defined a representation $V^{\lambda}$ of $G^{\lambda}$ in §2.1. We can see $T$ as a maximal torus of $G^{\lambda}$. If $\mu\in X_{*}(T)$, we have $(V^{\lambda})^{\mu}=V^{\lambda}\cap V^{\mu}$, $(G^{\lambda})^{\mu}=G^{\lambda}\cap G^{\mu}$, $(V^{\lambda})^{\mu\geqslant 0}$ and $P_{\mu,\lambda}\coloneqq P_{\mu}\cap G^{\lambda}$.
In this case, the induction diagram is
[…]
The definition of the induction map associated with this datum is
@@
If $\mu\preceq\lambda$, then $(V^{\lambda})^{\mu}=V^{\mu}$ and $(G^{\lambda})^{\mu}=G^{\mu}$, so that the induction $\mathrm{Ind}_{\mu,\lambda}$ is
@@
We obtain $\nu$ by composing a general one-parameter subgroup $\mathbf{G}_{\mathrm{m}}\rightarrow\mathbf{G}_{\mathrm{m}}^{\times 2}$, $t\mapsto(t^{a},t^{b})$ with $\lambda\times\mu$. This guarantees the first two equalities of the Lemma. To guarantee the last two, it suffices to take a general one-parameter subgroup with the constraint $b>0$. ∎
@@ 2.4. Associativity
Let $\lambda,\mu,\nu\in\mathrm{Hom}(\mathbf{G}_{\mathrm{m}},T)$. Then,
@@
By Lemma 2.6, we may assume that $\lambda\preceq\mu\preceq\nu$. This assumption only simplifies the notations in diagram (2.4). The equality then classically follows by base-change in the commutative diagram
[…]
@@ 2.5. Explicit formula for the parabolic induction
For $\lambda\in X_{*}(T)$ we define the induction kernel
$k_{\lambda}\coloneqq\frac{\prod_{\alpha\in\mathcal{W}^{\lambda>0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\lambda>0}(\mathfrak{g})}\alpha}\in\mathrm{Frac}(\mathrm{H}^{*}(\mathrm{pt}/T)).$
@@
It suffices to prove that $W^{\lambda}$ send $\mathcal{W}^{\lambda>0}(V)\subseteq\mathcal{W}(W)$ to itself, and similarly for $\mathcal{W}^{\lambda>0}(\mathfrak{g})$. This comes from the invariance of the pairing between characters and cocharacters: for any $w\in W$, $\langle w\cdot\lambda,w\cdot\alpha\rangle=\langle\lambda,\alpha\rangle$ and if $w\in W^{\lambda}$, $w\cdot\lambda=\lambda$. ∎
@@
Let $\lambda\colon\mathbf{G}_{\mathrm{m}}\rightarrow T$ be a cocharacter and $f\in\mathrm{H}^{*}(V^{\lambda}/G^{\lambda})$. We have
@@
The second equality follows from the $W^{\lambda}$-invariance of both $f$ (since $f\in\mathrm{H}^{*}(V^{\lambda}/T)^{W^{\lambda}}$) and $k_{\lambda}$ (Lemma 2.8). The first equality is a computation of Euler class exactly as in the proof of [kontsevich2011cohomological, Theorem 2] in the case of quiver representations. ∎
@@
$k_{\lambda,\mu}\coloneqq\frac{\prod_{\alpha\in\mathcal{W}^{\lambda>0}(V^{\mu})}\alpha}{\prod_{\alpha\in\mathcal{W}^{\lambda>0}(\mathfrak{g}^{\mu})}\alpha}.$
Proposition 2.10 (Augmentation).
For $f\in\mathrm{H}^{*}((V^{\mu})^{\lambda}/(G^{\mu})^{\lambda})$, we have
@@
2.6. Parabolic induction for symmetric representations
@@
Let $V$ be a symmetric representation of a reductive group $G$ and $\lambda\in X_{*}(T)$ a cocharacter of the maximal torus of $T$. Then, the representation $V^{\lambda}$ of $G^{\lambda}$ is also symmetric.
@@
It is immediate to check that $\mathcal{W}^{\prime}(V^{\lambda})=\{\alpha\in\mathcal{W}^{\prime}(V)\mid\langle\lambda,\alpha\rangle=0\}$ and for any $\alpha\in\mathcal{W}^{\prime}(V^{\lambda})$, we can identify the weight spaces $V^{\lambda}_{\alpha}=V_{\alpha}$. Therefore, the symmetry of $V$ implies that of $V^{\lambda}$. ∎
Recall the order relation on $X_{*}(T)$ defined in the introduction
[…]
@@
We let $\mathscr{P}_{V}\coloneqq X_{*}(T)/\sim$ be the quotient of $X_{*}(T)$ by this equivalence relation. If $\lambda\in X_{*}(T)$, $\overline{\lambda}$ denotes its class in $\mathscr{P}_{V}$.
It has an induced order relation still denoted by $\preceq$:
@@
Let $V$ be a symmetric representation of a reductive group $G$. If $\lambda,\mu\in X_{*}(T)$ are such that $\lambda\sim\mu$, then #$\#\mathcal{W}^{\lambda>0}(V)=\#\mathcal{W}^{\mu>0}(V)=\#\mathcal{W}^{\lambda<0}(V)=\#\mathcal{W}^{\mu<0}(V)$ and #$\#\mathcal{W}^{\lambda>0}(\mathfrak{g})=\#\mathcal{W}^{\mu>0}(\mathfrak{g})=\#\mathcal{W}^{\lambda<0}(\mathfrak{g})=\#\mathcal{W}^{\mu<0}(\mathfrak{g})$.
@@
We prove the statement for $V$. The statement for $\mathfrak{g}$ is obtained by specialising $V=\mathfrak{g}$ since $\mathfrak{g}$ under the adjoint action of $G$ is a symmetric representation of $G$.
Since $V$ is symmetric, we have #$\#\mathcal{W}^{\lambda>0}(V)=\#\mathcal{W}^{\lambda<0}(V)$ and moreover, #$\#\mathcal{W}(V)=\#\mathcal{W}^{\lambda>0}(V)+\#\mathcal{W}^{\lambda<0}(V)+\mathcal{W}(V^{\lambda})$. Since we have the same equalities with $\lambda$ replaced by $\mu$ and $\mathcal{W}^{\lambda=0}(V)=\mathcal{W}^{\mu=0}(V)$, we obtain #$\#\mathcal{W}^{\lambda>0}(V)=\#\mathcal{W}^{\mu>0}(V)$. The other equalities follow. ∎
@@
If $\lambda,\mu\in X_{*}(T)$ are such that $\lambda\sim\mu$, then $k_{\lambda}=\pm k_{\mu}$.
@@
If $\lambda\sim\mu$, we have $\mathcal{W}^{\lambda>0}(V)\sqcup\mathcal{W}^{\lambda<0}(V)=\mathcal{W}^{>\mu}(V)\sqcup\mathcal{W}^{<\mu}(V)$ and $\mathcal{W}^{\lambda>0}(\mathfrak{g})\sqcup\mathcal{W}^{\lambda<0}(\mathfrak{g})=\mathcal{W}^{>\mu}(\mathfrak{g})\sqcup\mathcal{W}^{<\mu}(\mathfrak{g})$ and since $V$ and $\mathfrak{g}$ are symmetric representations of $G$, $\mathcal{W}^{\lambda>0}(V)=-\mathcal{W}^{\lambda<0}(V)$, $\mathcal{W}^{>\mu}(V)=-\mathcal{W}^{<\mu}(V)$ and similarly for $\mathfrak{g}$. Therefore,
$(-1)^{\#\mathcal{W}^{\lambda>0}(V)-\#\mathcal{W}^{\lambda>0}(\mathfrak{g})}k_{\lambda}^{2}=\frac{\prod_{\alpha\in\mathcal{W}^{\lambda>0}(V)\sqcup\mathcal{W}^{\lambda<0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\lambda>0}(\mathfrak{g})\sqcup\mathcal{W}^{\lambda<0}(\mathfrak{g})}\alpha}=\frac{\prod_{\alpha\in\mathcal{W}^{\mu>0}(V)\sqcup\mathcal{W}^{\mu<0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu>0}(\mathfrak{g})\sqcup\mathcal{W}^{\mu<0}(\mathfrak{g})}\alpha}=(-1)^{\#\mathcal{W}^{\mu>0}(V)-\#\mathcal{W}^{\mu>0}(\mathfrak{g})}k_{\mu}^{2}.$
Since by Lemma 2.12 we have #$\#\mathcal{W}^{\lambda>0}(V)=\#\mathcal{W}^{\lambda<0}(V)=\#\mathcal{W}^{\mu>0}(V)=\#\mathcal{W}^{\mu<0}(V)$ and #$\#\mathcal{W}^{\lambda>0}(\mathfrak{g})=\#\mathcal{W}^{\lambda<0}(\mathfrak{g})=\#\mathcal{W}^{\mu>0}(\mathfrak{g})=\#\mathcal{W}^{\mu<0}(\mathfrak{g})$, we have $k_{\lambda}^{2}=k_{\mu}^{2}$ and therefore, $k_{\lambda}=\pm k_{\mu}$. ∎
@@
We assume that $V$ is a symmetric representation of $G$. Then, for any $\lambda,\mu\in X_{*}(T)$ such that $\overline{\lambda}=\overline{\mu}$, the induction maps $\mathrm{Ind}_{\lambda}$ and $\mathrm{Ind}_{\mu}$ differ by a sign. In particular, the images of $\mathrm{Ind}_{\lambda}$ and $\mathrm{Ind}_{\mu}$ inside $\mathrm{H}^{*+d}(V/G)$ coincide.
@@
Moreover, $G^{\lambda}=G^{\mu}$, $W^{\lambda}=W^{\mu}$ and by Lemma 2.13, $k_{\mu}=\pm k_{\lambda}$. This concludes. ∎
@@
By Lemma 2.14, we may assume that for some $w\in W$, $\mu=w\cdot\lambda$. Then, $\mathcal{W}^{\mu>0}(V)=w\cdot\mathcal{W}^{\lambda>0}(V)$ and $\mathcal{W}^{\mu>0}(\mathfrak{g})=w\cdot\mathcal{W}^{\lambda>0}(\mathfrak{g})$. Therefore, $w(k_{\lambda})=k_{\mu}$. Using the explicit formula for the induction product (Proposition 2.9), we obtain
@@
3. Tautological classes

### v6

@@ 2.2. Parabolic induction
For $\lambda\in X_{*}(T)$, we let $\mathbf{Q}_{V^{\lambda}/G^{\lambda}}^{\mathrm{vir}}\coloneqq\mathbf{Q}_{V^{\lambda}/G^{\lambda}}\otimes\mathscr{L}^{-d_{\lambda}/2}=\mathcal{IC}(V^{\lambda}/G^{\lambda})$ be the intersection complex monodromic mixed Hodge module of the smooth stack $V^{\lambda}/G^{\lambda}$. We refer to [DM20] for the necessary background regarding monodromic mixed Hodge modules, and in particular regarding the square-root $\mathscr{L}^{1/2}$ of the Tate twist $\mathscr{L}$. The Tate twist $\mathscr{L}\coloneqq\mathrm{H}^{*}_{\mathrm{c}}(\mathbf{A}^{1},\mathbf{Q})$ is a pure weight 0 complex of mixed Hodge structures, concentrated in degree $2$. Alternatively, for the purposes of this paper, the reader may prefer to consider instead constructible sheaves. The Tate twist -$\otimes\mathscr{L}$ is then replaced by the shift [2][-2].
@@
Moreover, the map $p_{\lambda}$ is proper and representable (Lemma 2.1), and so by dualizing the adjunction map
@@
by composing $(\imath_{\lambda})_{*}(\pi_{\lambda})_{*}$ applied to (2.2) with $\pi_{*}$ applied to (2.3) twisted by $\mathscr{L}^{r_{\lambda}+d_{\lambda}/2}$. We used the commutativity of (2.1) to identify $\pi_{*}(p_{\lambda})_{*}\mathbb{D}\mathbf{Q}_{V^{\lambda\geqslant 0}/P_{\lambda}}$ and $(\imath_{\lambda})_{*}(\pi_{\lambda})_{*}(q_{\lambda})_{*}\mathbb{D}\mathbf{Q}_{V^{\lambda\geqslant 0}/P_{\lambda}}$.
@@
Let $V$ be a weakly symmetric representation of $G$. Then, the induction morphism $\mathrm{Ind}_{\lambda}\colon\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/G^{\lambda})\rightarrow\mathrm{H}^{*+d}(V/G)$ preserves shifted cohomological degrees, that is $\mathrm{Ind}_{\lambda}(\mathrm{H}^{i+d_{\lambda}}(V^{\lambda}/G^{\lambda}))\subseteq\mathrm{H}^{i+d}(V/G)$ for any $i\in\mathbf{Z}$.
@@
The induction at the level of cohomology vector spaces is given by applying the derived global section functor to the morphism of complexes (2.4) (the sheafified induction). Moreover, by Lemma 2.6, $d/2=r_{\lambda}+d_{\lambda}/2$. Therefore, in the weakly symmetric case, the right-hand-side of (2.4) is precisely $\pi_{*}\mathbb{D}\mathbf{Q}_{V/G}^{\mathrm{vir}}$. Moreover, $\mathbb{D}\mathbf{Q}_{V/G}^{\mathrm{vir}}\cong\mathbf{Q}_{V/G}^{\mathrm{vir}}$ and $\mathbb{D}\mathbf{Q}_{V^{\lambda}/G^{\lambda}}^{\mathrm{vir}}\cong\mathbf{Q}_{V^{\lambda}/G^{\lambda}}^{\mathrm{vir}}$ since $V/G$ and $V^{\lambda}/G^{\lambda}$ are smooth and the chosen shifts/Tate twists (by the dimension) make the considered constant sheaves Verdier self-dual. ∎
@@ 2.3. Augmentation
Let $\lambda\in X_{*}(T)$. We defined a representation $V^{\lambda}$ of $G^{\lambda}$ in §2.1. The induction morphisms, which we have defined for $(V,G)$, can also be defined in the exact same way for $(V^{\lambda},G^{\lambda})$. We can see $T$ as a maximal torus of $G^{\lambda}$. If $\mu\in X_{*}(T)$, we have $(V^{\lambda})^{\mu}=V^{\lambda}\cap V^{\mu}$, $(G^{\lambda})^{\mu}=G^{\lambda}\cap G^{\mu}$, $(V^{\lambda})^{\mu\geqslant 0}=V^{\lambda}\cap V^{\mu\geqslant 0}$ and $P_{\mu,\lambda}\coloneqq P_{\mu}\cap G^{\lambda}$.
In this case, the induction diagram (2.1) becomes
$(V^{\lambda})^{\mu\geqslant 0}/P_{\mu,\lambda}$$(V^{\lambda})^{\mu}/(G^{\lambda})^{\mu}$$V^{\lambda}/G^{\lambda}$$(V^{\lambda})^{\mu}/\!\!/(G^{\lambda})^{\mu}$$V^{\lambda}/\!\!/G^{\lambda}$$\scriptstyle{q_{\lambda,\mu}}$$\scriptstyle{p_{\lambda,\mu}}$$\scriptstyle{\pi_{\lambda,\mu}}$$\scriptstyle{\pi_{\lambda}}$$\scriptstyle{\imath_{\lambda,\mu}}$
The definition of the induction map associated with this datum following §2.2 is
@@
If $\mu\preceq\lambda$ (1.2), then $(V^{\lambda})^{\mu}=V^{\mu}$ and $(G^{\lambda})^{\mu}=G^{\mu}$, so that the induction $\mathrm{Ind}_{\mu,\lambda}$ is
@@
We obtain $\nu$ by composing a general one-parameter subgroup $\zeta\colon\mathbf{G}_{\mathrm{m}}\rightarrow\mathbf{G}_{\mathrm{m}}^{\times 2}$, $t\mapsto(t^{a},t^{b})$ with $\lambda\times\mu$. This guarantees the first two equalities of the Lemma. To guarantee the last two, it suffices to take a general one-parameter subgroup with the constraint $b>0$ by replacing $\zeta$ by $\zeta^{-1}$ if necessary. ∎
@@ 2.4. Associativity
Let $\lambda,\mu,\nu\in X_{*}(T)$ be such that $V^{\lambda}\subseteq V^{\mu}\subseteq V^{\nu}$, $G^{\lambda}\subseteq G^{\mu}\subseteq G^{\nu}$ and $(V^{\nu})^{\mu\geqslant 0}\subseteq(V^{\nu})^{\lambda\geqslant 0}$, $P_{\mu,\nu}\subseteq P_{\lambda,\nu}$. Then,
@@
The equality classically follows by base-change in the commutative diagram
(2.5) $(V^{\nu})^{\lambda\geqslant 0}/P_{\lambda,\nu}$$(V^{\mu})^{\lambda\geqslant 0}/P_{\lambda,\mu}$$(V^{\nu})^{\mu\geqslant 0}/P_{\mu,\nu}$$V^{\lambda}/G^{\lambda}$$V^{\mu}/G^{\mu}$$V^{\nu}/G^{\nu}$$\scriptstyle{q_{\nu,\lambda}}\scriptstyle{\lrcorner}$$\scriptstyle{p_{\nu,\lambda}}$$\scriptstyle{q_{\mu,\lambda}}$$\scriptstyle{p_{\mu,\lambda}}$$\scriptstyle{q_{\nu,\mu}}$$\scriptstyle{p_{\nu,\mu}}$
@@ 2.5. Explicit formula for the parabolic induction
If $\mathfrak{t}_{\mathbf{Q}}$ is a rational form of the Lie algebra of $T$, we have $\mathrm{H}^{*}_{T}(\mathrm{pt})\cong\mathrm{Sym}(\mathfrak{t}^{*})$. Any character $\alpha\in X_{*}(T)$ may be identified with an element of $\mathfrak{t}^{*}$ and hence $\mathrm{H}^{*}_{T}(\mathrm{pt})$ via $\alpha\mapsto\mathrm{d}\alpha(1)$ (see §1.4). By abuse, we will still denote by $\alpha$ the corresponding element of $\mathrm{H}^{*}_{T}(\mathrm{pt})$.
For $\lambda\in X_{*}(T)$, we define the induction kernel
$k_{\lambda}\coloneqq\frac{\prod_{\alpha\in\mathcal{W}^{\lambda<0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\lambda<0}(\mathfrak{g})}\alpha}\in\mathrm{Frac}(\mathrm{H}^{*}(\mathrm{pt}/T))\,.$
@@
It suffices to prove that $W^{\lambda}$ sends $\mathcal{W}^{\lambda<0}(V)\subseteq\mathcal{W}(V)$ to itself, and similarly for $\mathcal{W}^{\lambda<0}(\mathfrak{g})$. This comes from the invariance of the pairing between characters and cocharacters: for any $w\in W$, $\langle w\cdot\lambda,w\cdot\alpha\rangle=\langle\lambda,\alpha\rangle$ and if $w\in W^{\lambda}$, $w\cdot\lambda=\lambda$. Therefore, for any $\alpha\in X^{*}(T)$, $\langle\lambda,\alpha\rangle<0$ if and only if $\langle\lambda,w\cdot\alpha\rangle<0$. ∎
@@
Let $\lambda\colon\mathbf{G}_{\mathrm{m}}\rightarrow T$ be a cocharacter and $f\in\mathrm{H}^{*}(V^{\lambda}/G^{\lambda})\cong\mathrm{H}^{*}(\mathrm{pt}/T)^{W^{\lambda}}$. We have
@@
The second equality follows from the $W^{\lambda}$-invariance of both $f$ (since $f\in\mathrm{H}^{*}(V^{\lambda}/G^{\lambda})=\mathrm{H}^{*}(V^{\lambda}/T)^{W^{\lambda}}$) and $k_{\lambda}$ (Lemma 2.10). The first equality is a computation of Euler classes exactly as in the proof of [KS11, Theorem 2] in the case of quiver representations. ∎
@@
$k_{\lambda,\mu}\coloneqq\frac{\prod_{\alpha\in\mathcal{W}^{\lambda<0}(V^{\mu})}\alpha}{\prod_{\alpha\in\mathcal{W}^{\lambda<0}(\mathfrak{g}^{\mu})}\alpha}\in\mathrm{Frac}(\mathrm{H}^{*}_{T}(\mathrm{pt}))\,.$
Proposition 2.12 (Augmentation of Proposition 2.11).
For $f\in\mathrm{H}^{*}((V^{\mu})^{\lambda}/(G^{\mu})^{\lambda})\cong\mathrm{H}^{*}(\mathrm{pt}/T)^{W^{\lambda}\cap W^{\mu}}$, we have
@@
2.6. Parabolic induction for weakly symmetric representations
@@
Let $V$ be a (weakly) symmetric representation of a reductive group $G$ and $\lambda\in X_{*}(T)$ a cocharacter of the maximal torus of $T$. Then, the representation $V^{\lambda}$ of $G^{\lambda}$ is also (weakly) symmetric.
@@
It is immediate to check that $\mathcal{W}^{\prime}(V^{\lambda})=\{\alpha\in\mathcal{W}^{\prime}(V)\mid\langle\lambda,\alpha\rangle=0\}$ and for any $\alpha\in\mathcal{W}^{\prime}(V^{\lambda})$, we can identify the weight spaces $V^{\lambda}_{\alpha}=V_{\alpha}$. Therefore, the (weak) symmetry of $V$ implies that of $V^{\lambda}$. ∎
Recall the order relation on $X_{*}(T)$ defined in the introduction (1.2)
$\lambda\preceq\mu\overset{\text{def.}}{\iff}\left\{\begin{aligned} V^{\lambda}\subseteq V^{\mu}\\ \mathfrak{g}^{\lambda}\subseteq\mathfrak{g}^{\mu}\,.\end{aligned}\right$.
@@
We let $\mathscr{P}_{V}\coloneqq X_{*}(T)/\sim$ be the quotient of $X_{*}(T)$ by this equivalence relation. If $\lambda\in X_{*}(T)$, $\overline{\lambda}$ denotes its class in $\mathscr{P}_{V}$. It has an induced order relation still denoted by $\preceq$:
@@
Let $V$ be a weakly symmetric representation of a reductive group $G$. If $\lambda,\mu\in X_{*}(T)$ are such that $\lambda\sim\mu$, then
  1. (1)
    $\mathcal{W}^{\lambda>0}(V)=\#\mathcal{W}^{\mu>0}(V)=\#\mathcal{W}^{\lambda<0}(V)=\#\mathcal{W}^{\mu<0}(V)$,
  2. (2)
    $\mathcal{W}^{\lambda>0}(\mathfrak{g})=\#\mathcal{W}^{\mu>0}(\mathfrak{g})=\#\mathcal{W}^{\lambda<0}(\mathfrak{g})=\#\mathcal{W}^{\mu<0}(\mathfrak{g})$.
@@
We prove the statement for $V$. The statement for $\mathfrak{g}$ is obtained by specialising $V=\mathfrak{g}$ since $\mathfrak{g}$ under the adjoint action of $G$ is a symmetric representation of $G$ (by the existence of invariant pairings, nondegenerate up to the center of $\mathfrak{g}$, Example 2.4).
Since $V$ is weakly symmetric, we have $\mathcal{W}^{\lambda>0}(V)=\#\mathcal{W}^{\lambda<0}(V)$ and moreover, $\mathcal{W}(V)=\#\mathcal{W}^{\lambda>0}(V)+\#\mathcal{W}^{\lambda<0}(V)+\mathcal{W}(V^{\lambda})$. Since we have the same equalities with $\lambda$ replaced by $\mu$ and $\mathcal{W}^{\lambda=0}(V)=\mathcal{W}^{\mu=0}(V)$, we obtain $\mathcal{W}^{\lambda>0}(V)=\#\mathcal{W}^{\mu>0}(V)$. The other equalities follow. ∎
@@
If $\lambda,\mu\in X_{*}(T)$ are such that $\lambda\sim\mu$, then $k_{\lambda}=\xi k_{\mu}$ for some $\xi\in\mathbf{Q}^{*}$.
@@
If $\lambda\sim\mu$, we have $\mathcal{W}^{\lambda>0}(V)\sqcup\mathcal{W}^{\lambda<0}(V)=\mathcal{W}^{\mu>0}(V)\sqcup\mathcal{W}^{\mu<0}(V)$ and $\mathcal{W}^{\lambda>0}(\mathfrak{g})\sqcup\mathcal{W}^{\lambda<0}(\mathfrak{g})=\mathcal{W}^{\mu>0}(\mathfrak{g})\sqcup\mathcal{W}^{\mu<0}(\mathfrak{g})$ and since $V$ and $\mathfrak{g}$ are weakly symmetric representations of $G$, $\overline{\mathcal{W}^{\lambda>0}(V)}=-\overline{\mathcal{W}^{\lambda<0}(V)}$, $\overline{\mathcal{W}^{\mu>0}(V)}=-\overline{\mathcal{W}^{\mu<0}(V)}$ and similarly for $\mathfrak{g}$. Therefore, for any $\alpha\in\mathcal{W}^{\lambda<0}(V)$ (resp. $\mathcal{W}^{\lambda<0}(\mathfrak{g})$), there exists $\xi_{\alpha}\in\mathbf{Q}^{*}$ (resp. $\zeta_{\alpha}\in\mathbf{Q}^{*}$) such that
$\begin{matrix}\mathcal{W}^{\lambda<0}(V)&\rightarrow&\mathcal{W}^{\mu<0}(V)\\ \alpha&\mapsto&\xi_{\alpha}\alpha\end{matrix}$
and
$\begin{matrix}\mathcal{W}^{\lambda<0}(\mathfrak{g})&\rightarrow&\mathcal{W}^{\mu<0}(\mathfrak{g})\\ \alpha&\mapsto&\zeta_{\alpha}\alpha\end{matrix}$
are bijections. Then,
$\frac{\prod_{\alpha\in\mathcal{W}^{\lambda<0}(V)}\xi_{\alpha}}{\prod_{\alpha\in\mathcal{W}^{\lambda<0}(\mathfrak{g})}\zeta_{\alpha}}k_{\lambda}=k_{\mu}\,.$
We may therefore set $\xi\coloneqq\frac{\prod_{\alpha\in\mathcal{W}^{\lambda<0}(\mathfrak{g})}\zeta_{\alpha}}{\prod_{\alpha\in\mathcal{W}^{\lambda<0}(V)}\xi_{\alpha}}$. ∎
@@
We assume that $V$ is a weakly symmetric representation of $G$. Then, for any $\lambda,\mu\in X_{*}(T)$ such that $\overline{\lambda}=\overline{\mu}$, the induction maps $\mathrm{Ind}_{\lambda}$ and $\mathrm{Ind}_{\mu}$ differ by a non-zero scalar. In particular, the images of $\mathrm{Ind}_{\lambda}$ and $\mathrm{Ind}_{\mu}$ inside $\mathrm{H}^{*+d}(V/G)$ coincide.
@@
Moreover, $G^{\lambda}=G^{\mu}$, $W^{\lambda}=W^{\mu}$ and by Lemma 2.15, $k_{\lambda}=\xi k_{\mu}$. This concludes the proof. ∎
@@
By Lemma 2.16, we may assume that for some $w\in W$, $\mu=w\cdot\lambda$. Then, $\mathcal{W}^{\mu<0}(V)=w\cdot\mathcal{W}^{\lambda<0}(V)$ and $\mathcal{W}^{\mu<0}(\mathfrak{g})=w\cdot\mathcal{W}^{\lambda<0}(\mathfrak{g})$. Therefore, $w(k_{\lambda})=k_{\mu}$. Using the explicit formula for the induction product (Proposition 2.11), we obtain
@@
3. Action of tautological classes on the cohomology of the stack